PROBLEM C
Attention : Please answer all problems in PROBLEM C with the right and short explanation, explain by using the formula (if needed).
A. A firm can have a high trailing PIE ratio, yet have a low expected earnings growth rate in
the future. Is this so ?
B. For a firm with a normal trailing PIE ratio, expected future residual earnings must be the
same as current residual earnings. Correct ?
C. Does an increase in financial leverage increase or decrease the (levered) P/E ratio ?
D. Wall Street analyst predicted in 2009 (1 year after crisis 2008) that, after the considerable
deleveraging during the financial crisis, firms would begin once again to lever up with
more borrowing. "They must defend their return on equity," he claimed. As a result,
"investors should look for a rise in dividends and share buy-backs and an expansion of PIE
multiples, leading to equity market outperformance." Is he correct ?
E. The higher the anticipated return on net operating assets (RNOA) relative to the anticipated
growth in net operating assets, the higher will be the unlevered price to book ratio. Is this
correct ?
Answer #1

A) Yes, a firm can have high trailing PE ratio yet have a low expected earning growth. It can be understand by this formula.

Trailing PE Ratio = Current Market price devided by last 12 months Earning per share

Assume Current market price = 4000 Rs

last 12 month EPS total = 100 Rs

so Trainling PE = 4000/100 = 40 times

Now as the future EPS gets increases it will impact the PE ratio to reduce

for example -

Assume current market price = 4000 Rs

forecasted Future EPS due to change in the future prospect of business = 200 Rs

Now Future PE ratio = 4000/ 200 = 20 time

so to conclude as the EPS increases the PE ratio decreases which tells us the low future Earnings

B) Not Really. The future earnings of a company depends upon its future business prospects which is changable in nature

C) Financial leverage comes into the picture when company finances its most of the assets through borrowings. and more borrowing attracts more interest payment which decreases the earnings of the company.

so when the earnings are less the EPS of the company would be less which in returen increase the PE multiple and create the situation of overvaluation of stock .

D) Yes the analyst is right as the companies after the crisis will more depend upon the borrowings the interest payment on the borrowings will also be increased. but due to share buyback companies will try to project a positive scenario in the market with increased EPS as after the buyback the no. of outstanding shares will decrease and in result the Demand of that stock will rise and people will try to buy more and more which will increase the market price of the comapnies

E) A higher return on Net operationg asset indicates that the company is utilizing its assets very well in order to generate income for the company. if the company having unlevered book value means there is no leverage or its already been paid . so as a result levered price to book value will get reduced as there is no leverage so total assets will be high in denominator which will make the ratio lower